# Representing Molecules as Random Walks Over Interpretable Grammars

### v1

@@ 3.3.3 Molecular Generation
To generate a molecule $M$, we apply the learned grammar forward to sample edges to traverse during the random walk, as depicted in Fig. 2. Each sampled edge either attaches a new motif to the current $M$, or backtracks to a previous motif. Details on the algorithm are given in Appendix G.
@@ 4 Results & Analysis
Our experiments quantitatively answer the following questions: 1) How well does our method perform on property prediction for our setting of interest? 2) How well does our representation work for the generation of novel molecules? We also include three ablations to answer: 3) How important are expert motifs, compared to heuristic-based motifs? 4) How data-efficient and runtime-efficient is our method? 5) How does our method compare with a simple baseline that only consider motifs? Our qualitative analysis answers the following questions: 6) How can our model extract rules and are they meaningful to an expert? 7) How interpretable are our model’s learnt representations?
@@ 4.1 Datasets and Baselines
Baselines.
To address question 1), we compare with pretrained GNNs (PN (Stanley et al., 2021) and Pre-trained GIN (Hu et al., 2020)), a specialized GNN model for property prediction (wD-MPNN (Aldeghi & Coley, 2022)), two SOTA pretrained models for molecular representation learning (MolCLR (Wang et al., 2022) and UniMol (Zhou et al., 2023)) and two SOTA subgraph-based methods (ESAN (Bevilacqua et al., 2022) and HM-GNN (Shui & Karypis, 2020)). To address question 2), we compare with both Geo-DEG, the SOTA on small dataset property prediction, and its generative variant, DEG, for molecular generation.
@@ 4.2 Results
We report the mean absolute error (MAE) and coefficient of determination ($R^{2}$) over normalized prediction values for GC and HOPV, and the accuracy and AUC for PTC. For each (dataset, property) pair, we perform an 80-20 train-test split over 3 random seeds and report the mean and standard deviation. For molecular generation, we report commonly used metrics (Polykovskiy et al., 2020; Guo et al., 2023a) [^5] We drop the metric of Membership, which tests whether certain group(s) characteristic of membership to the chemical class are present. Our method, by design, combines chemically meaningful groups, so it would be pointless to test for it.: Validity/Uniqueness: Percentage of chemically valid/unique molecules; Diversity: Average pairwise molecular distance among generated molecules; Retro* Score (RS): Success rate of Retro* (Chen et al., 2020) which was trained to find a retrosynthesis path to build a molecule from a list of commercially available ones.
@@ 4.2.1 Property prediction
Group
MAE $\downarrow$
$0.47\pm 0.09$
$0.51\pm 0.06$
$0.34\pm 0.12$
$0.76\pm 0.30$
$0.68\pm 0.05$
$0.26\pm 0.10$
$0.38\pm 0.13$
$0.26\pm 0.11$
$\mathbf{0.25\pm 0.09}$
$0.27\pm 0.08$
$\mathbf{R^{2}}\uparrow$
$0.41\pm 0.12$
$-0.39\pm 0.62$
$0.56\pm 0.20$
7.56±7.71-7.56$\pm$-7.71- 7.56 ± - 7.71
$0.19\pm 0.09$
$0.68\pm 0.20$
$0.47\pm 0.25$
$0.70\pm 0.20$
$\mathbf{0.80\pm 0.15}$
$0.74\pm 0.15$
HOPV
MAE $\downarrow$
$0.36\pm 0.03$
$0.37\pm 0.02$
$0.40\pm 0.02$
$0.42\pm 0.02$
$0.38\pm 0.02$
$0.34\pm 0.03$
$0.31\pm 0.03$
$0.30\pm 0.02$ $0.30\pm 0.05$
$\mathbf{0.22\pm 0.15}$
$\mathbf{R^{2}}\uparrow$
$0.69\pm 0.04$
$0.66\pm 0.06$
$0.65\pm 0.05$
$0.65\pm 0.04$
$0.66\pm 0.03$
$0.68\pm 0.03$
$0.70\pm 0.02$
$0.74\pm 0.03$
$\mathbf{0.80\pm 0.06}$
$0.77\pm 0.12$
PTC
Acc $\uparrow$
$0.67\pm 0.06$
$0.64\pm 0.08$
$0.66\pm 0.07$
$0.61\pm 0.08$
$0.62\pm 0.09$
$0.60\pm 0.03$
$0.57\pm 0.05$
$0.69\pm 0.07$
$\mathbf{0.70\pm 0.01}$
$0.67\pm 0.02$
AUC $\uparrow$
$0.70\pm 0.05$
$0.68\pm 0.06$
$0.69\pm 0.06$
$0.65\pm 0.07$
$0.66\pm 0.07$
$0.66\pm 0.05$
$0.67\pm 0.06$
$\mathbf{0.71\pm 0.07}$
$0.69\pm 0.03$
$0.66\pm 0.05$
Our method, with expert motifs, achieves the best $R^{2}$ by a wide margin of $0.10$ and $0.06$ over the second best method on regression datasets Group Contribution and HOPV and the highest accuracy on PTC. With heuristic motifs, our method remains competitive to Geo-DEG, achieving higher $R^{2}$ on both regression datasets and accuracy within standard deviation on PTC. Interestingly, using heuristic-based motifs in HOPV, achieves significantly (27%) lower MAE than expert-based motifs and Geo-DEG. To answer question 3), we see that the ablation suggests expert motifs are generally better, but may be more sensitive to outliers than heuristic-based motifs. We observe experts are generally better at identifying special cases that heuristics are unaware of, but heuristics are more consistent. This reflects how $R^{2}$ is generally more sensitive to outliers than MAE. We describe our experts’ annotation criteria in Appendix Section A and do an in-depth case study on HOPV in Appendix Section G.
@@ 4.2.2 Molecular generation
To answer question 2), we see in Table 3 that our method produces comparably more diverse molecules than the training dataset (+0.03 on HOPV, -0.01 on PTC) and significantly more synthesizable molecules than the previous SOTA, DEG (+43% on both datasets). On HOPV, our retrosynthetic planner finds synthesis paths at a 14% higher rate on our novel molecules than the original dataset, a careful collation of experimental photovoltaic designs (Lopez et al., 2016). This is encouraging to experimentalists whose work is contingent on the designs’ feasibility for synthesis.
Table 3: Results on molecular generation for HOPV (top) and PTC (bottom); for both datasets, we generate 1000 novel molecules.
Datasets Methods Valid Unique Diversity RS
HOPV Train Data 100% 100% 0.86 41%
DEG 100% 100% 0.94 12%
Ours 100% 100% 0.89 58%
PTC Train Data 100% 100% 0.94 87%
DEG 100% 100% 0.94 17%
Ours 100% 100% 0.93 60%
@@ 4.3.1 Ablation: vary training dataset size
Table 4: We fix the same testing data over each dataset.
Ablation/Dataset
HOPV PTC Group Contribution
Train MAE
Train $R^{2}$
Test MAE
Test $R^{2}$
Train MAE
Train $R^{2}$
Test MAE
Test $R^{2}$
Train MAE
Train $R^{2}$
Test MAE
Test $R^{2}$
Bag-of-Motifs
$\mathbf{0.011\pm}$ $\mathbf{0.004}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.521\pm$ $0.031$ $0.446\pm$ $0.125$ $\mathbf{0.996\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $\mathbf{0.581\pm}$ $\mathbf{0.018}$ $\mathbf{0.612\pm}$ $\mathbf{0.029}$ $\mathbf{0.000\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.493\pm$ $0.143$ $0.214\pm$ $0.404$
Bag-of-Motifs (-expert)
$0.014\pm$ $0.002$ $0.997\pm$ $0.060$ $\mathbf{0.486\pm}$ $\mathbf{0.025}$ $\mathbf{0.489\pm}$ $\mathbf{0.062}$ $\mathbf{0.996\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.529\pm$ $0.031$ $0.609\pm$ $0.031$ $\mathbf{0.000\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $\mathbf{0.481\pm}$ $\mathbf{0.174}$ $\mathbf{0.257\pm}$ $\mathbf{0.453}$
4.3.2 Ablation of Bag-of-Motifs
@@
To answer question 5), we ablate our hierarchical information and retain only the motif co-occurrence information. For each molecule, we obtain a feature vector that concatenates a) the occurrence counts of all motifs and b) the Morgan fingerprint of the molecule. We train an XGBoost regressor/classifier on top of these features (details in Appendix E). As shown in Table 4, this baseline has enough capacity to overfit the training data but fails to generalize. This allows us to conclude in the absence of a proper representation, motif occurrence information is not sufficient for generalization. Interestingly, expert-level motifs are not superior to heuristic-based motifs in this featurization. This suggests that the quality of motifs are not relevant in the absence of a hierarchical representation that incorporates the fine-grained features of each individual motif.
@@ 4.4.1 Rule extraction from grammar
To answer question 6), we extract context-sensitive grammar rules from our trained model. We perform best-first search over random walk trajectories, beginning with base trajectories corresponding to each group $v\in G$. We only expand trajectories with transition probabilities above a minimum threshold. Each trajectory that reaches a transition with probability of 1 is extracted as a “hard” context-sensitive rule. We depict two such rules in Figure 6, with a more exhaustive compilation in Appendix Section F.

### v2

@@ 3.3.3 Molecular Generation
To generate a molecule $M$, we apply the learned grammar forward to sample edges to traverse during the random walk, as depicted in Fig. 2. Each sampled edge either attaches a new motif to the current $M$, or backtracks to a previous motif. Details on the algorithm are given in Appendix F.
@@ 4 Results & Analysis
Our experiments quantitatively answer the following questions: 1) How well does our method perform on property prediction for our setting of interest? 2) How well does our representation work for the generation of novel molecules, compared with both SOTA symbolic and deep molecular generative models? We also include three ablations to answer: 3) How important are expert motifs, compared to heuristic-based motifs? 4) How data-efficient and runtime-efficient is our method? 5) How does our method compare with other motif-based predictors? Our qualitative analysis answers the following questions: 6) How interpretable is our learnt grammar to an expert? 7) How can we analyze the model’s learnt representations?
@@ 4.1 Datasets and Baselines
Baselines. To address question 1), we compare with pretrained GNNs (PN (Stanley et al., 2021) and Pre-trained GIN (Hu et al., 2020)), a specialized GNN model for property prediction (wD-MPNN (Aldeghi & Coley, 2022)), two SOTA pretrained models for molecular representation learning (MolCLR (Wang et al., 2022) and UniMol (Zhou et al., 2023)) and two SOTA subgraph-based methods (ESAN (Bevilacqua et al., 2022) and HM-GNN (Shui & Karypis, 2020)). To address question 2), we compare with both Geo-DEG, the SOTA on small dataset property prediction, and its generative variant, DEG, for molecular generation.
@@ 4.2 Results
We report the mean absolute error (MAE) and coefficient of determination ($R^{2}$) over normalized prediction values for GC and HOPV, and the accuracy and AUC for PTC. For each (dataset, property) pair, we perform an 80-20 train-test split over 3 random seeds and report the mean and standard deviation. For molecular generation, we report commonly used metrics (Polykovskiy et al., 2020; Guo et al., 2023a): Validity/Uniqueness/Novelty: Percentage of chemically valid/unique/novel molecules; Diversity: Average pairwise molecular distance among generated molecules; Retro* Score (RS): Success rate of Retro* (Chen et al., 2020) which was trained to find a retrosynthesis path to build a molecule from a list of commercially available ones. We add the metric of Membership, which tests whether certain motif(s) characteristic of membership to the chemical class are present, primarily as a sanity check. Our method, by design, can achieve 100% if the random walk initializes at the characteristic motif. See A.1 for further discussion.
@@ 4.2.1 Property prediction
To answer question 1), we see in Table 2 that our method, with expert motifs, achieves the best $R^{2}$ by a wide margin of $0.10$ and $0.06$ over the second best method on regression datasets GC and HOPV and the highest accuracy on PTC. With heuristic motifs, our method remains competitive to Geo-DEG, achieving higher $R^{2}$ on both regression datasets and accuracy within standard deviation on PTC. Interestingly, using heuristic-based motifs in HOPV, achieves significantly (27%) lower MAE than expert-based motifs and Geo-DEG. To answer question 3), we see that the ablation suggests expert motifs are generally better, but may be more sensitive to outliers than heuristic-based motifs. We observe experts are generally better at identifying special cases that heuristics are unaware of, but heuristics are more consistent. This reflects how $R^{2}$ is generally more sensitive to outliers than MAE. We describe our experts’ annotation criteria in Appendix A and do an in-depth case study on HOPV in Appendix G.
@@
Group MAE $\downarrow$ $0.47\pm 0.09$ $0.51\pm 0.06$ $0.34\pm 0.12$ $0.76\pm 0.30$ $0.68\pm 0.05$ $0.26\pm 0.10$ $0.38\pm 0.13$ $0.26\pm 0.11$ $\mathbf{0.25\pm 0.09}$ $0.27\pm 0.08$
$\mathbf{R^{2}}\uparrow$ $0.41\pm 0.12$ $-0.39\pm 0.62$ $0.56\pm 0.20$ 7.56±7.71-7.56$\pm$-7.71- 7.56 ± - 7.71 $0.19\pm 0.09$ $0.68\pm 0.20$ $0.47\pm 0.25$ $0.70\pm 0.20$ $\mathbf{0.80\pm 0.15}$ $0.74\pm 0.15$
HOPV MAE $\downarrow$ $0.36\pm 0.03$ $0.37\pm 0.02$ $0.40\pm 0.02$ $0.42\pm 0.02$ $0.38\pm 0.02$ $0.34\pm 0.03$ $0.31\pm 0.03$ $0.30\pm 0.02$ $0.30\pm 0.05$ $\mathbf{0.22\pm 0.15}$
$\mathbf{R^{2}}\uparrow$ $0.69\pm 0.04$ $0.66\pm 0.06$ $0.65\pm 0.05$ $0.65\pm 0.04$ $0.66\pm 0.03$ $0.68\pm 0.03$ $0.70\pm 0.02$ $0.74\pm 0.03$ $\mathbf{0.80\pm 0.06}$ $0.77\pm 0.12$
PTC Acc $\uparrow$ $0.67\pm 0.06$ $0.64\pm 0.08$ $0.66\pm 0.07$ $0.61\pm 0.08$ $0.62\pm 0.09$ $0.60\pm 0.03$ $0.57\pm 0.05$ $0.69\pm 0.07$ $\mathbf{0.70\pm 0.01}$ $0.67\pm 0.02$
AUC $\uparrow$ $0.70\pm 0.05$ $0.68\pm 0.06$ $0.69\pm 0.06$ $0.65\pm 0.07$ $0.66\pm 0.07$ $0.66\pm 0.05$ $0.67\pm 0.06$ $\mathbf{0.71\pm 0.07}$ $0.69\pm 0.03$ $0.66\pm 0.05$
@@ 4.2.2 Molecular generation
To answer question 2), we see in Table 3 that our method produces comparably more diverse molecules than the training dataset (+0.03 on HOPV, -0.01 on PTC) and significantly more synthesizable molecules than the previous SOTA, DEG (+39% on HOPV, +22% on PTC). On HOPV, our retrosynthetic planner finds synthesis paths at a 14% higher rate on our novel molecules than the original dataset, a careful collation of experimental photovoltaic designs (Lopez et al., 2016). This is encouraging to experimentalists whose work is contingent on the designs’ feasibility for synthesis. We also compare our methods with established VAE-based molecular generative models such as (Jin et al., 2018) and its follow-up work (Jin et al., 2020) which includes larger structural motifs. Furthermore, we modified the implementation of Hier-VAE to incorporate our epert motifs. For all three cases, we follow the default settings, train until convergence, and use the checkpoint with the lowest loss to sample 1000 molecules. We observe that both VAE-based methods struggle to generate sufficiently unique molecules, with only 11%-43% (HOPV) and 8%-28% (PTC) of the 1000 generated molecules being unique. This is despite sampling from a Gaussian noise distribution. Meanwhile, our model generates 100% unique and novel molecules, while ensuring a high internal diversity second only to DEG. For reference, (Jin et al., 2018, 2020) trained and evaluated their methods on  250K molecules extracted from ZINC (Sterling & Irwin, 2015) and a polymer dataset containing 86K polymers. Meanwhile, our datasets contain only  100-300 molecules and, in the case of HOPV, feature much larger molecules. Rather than using an encoder-decoder setup which requires significantly more data to learn the mapping to and from a latent space, our generative model explicitly captures the transition probabilities over traversing the symbolic space of structural motifs. Our grammar derivation can easily be conditioned by a set-based memory to apply a diverse set of transition rules. This leads to more unique, diverse, and, most importantly, synthesizable structures.
Table 3: Results on molecular generation for HOPV (top) and PTC (bottom); for both datasets, we generate 1000 novel molecules. Refer to Appendix A.1 for more details on Membership.
Datasets Methods Valid Unique Novel Diversity RS Memb.
HOPV Train Data 100% 100% N/A 0.86 51% 100%
DEG 100% 98% 99% 0.93 19% 46%
JT-VAE 100% 11% 100% 0.77 99% 84%
Hier-VAE 100% 43% 96% 0.87 79% 76%
Hier-VAE (+expert) 100% 29% 92% 0.86 84% 82%
Ours 100% 100% 100% 0.89 58% 71%
PTC Train Data 100% 100% N/A 0.94 87% 30%
DEG 100% 88% 87% 0.95 38% 27%
JT-VAE 100% 8% 80% 0.83 96% 27%
Hier-VAE 100% 20% 85% 0.91 92% 25%
Hier-VAE (+expert) 100% 28% 75% 0.93 90% 17%
Ours 100% 100% 100% 0.93 60% 22%
@@ 4.3.1 Ablation: vary training dataset size
Table 4: Ablation study on overfitting and generalization, vs other motif-based baselines, with and w/o expert motifs. Best result is bolded.
Ablation/Dataset HOPV PTC Group Contribution
Train MAE $\downarrow$ Train $R^{2}$uparrow Test MAE $\downarrow$ Test $R^{2}$uparrow Train Acc $\uparrow$ Train AUC $\uparrow$ Test Acc $\uparrow$ Test AUC $\uparrow$ Train MAE $\downarrow$ Train $R^{2}$uparrow Test MAE $\downarrow$ Test $R^{2}$uparrow
Bag-of-Motifs $0.014\pm$ $0.002$ $0.997\pm$ $0.001$ ${0.486\pm}$ 0.0250.025{0.025}0.025 ${0.489\pm}$ 0.0620.062{0.062}0.062 ${0.996\pm}$ 0.0000.000{0.000}0.000 $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.529\pm$ $0.031$ $0.609\pm$ $0.031$ $\mathbf{0.000\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ ${0.481\pm}$ 0.1740.174{0.174}0.174 ${0.257\pm}$ 0.4530.453{0.453}0.453
Bag-of-Motifs (+expert) $\mathbf{0.011\pm}$ $\mathbf{0.004}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.521\pm$ $0.031$ $0.446\pm$ $0.125$ ${0.996\pm}$ 0.0000.000{0.000}0.000 $\mathbf{1.000\pm}$ $\mathbf{0.000}$ ${0.581\pm}$ 0.0180.018{0.018}0.018 ${0.612\pm}$ 0.0290.029{0.029}0.029 $\mathbf{0.000\pm}$ $\mathbf{0.000}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ $0.493\pm$ $0.143$ $0.214\pm$ $0.404$
HM-GNN ${0.366\pm}$ 0.0350.035{0.035}0.035 ${0.686\pm}$ 0.0660.066{0.066}0.066 $0.473\pm$ $0.019$ $0.441\pm$ $0.065$ ${0.915\pm}$ 0.0330.033{0.033}0.033 ${0.966\pm}$ 0.0160.016{0.016}0.016 $\mathbf{0.710\pm}$ $\mathbf{0.023}$ ${{0.678\pm}}$ 0.0400.040{{0.040}}0.040 ${0.281\pm}$ 0.0640.064{0.064}0.064 ${0.717\pm}$ 0.1370.137{0.137}0.137 $0.362\pm$ $0.113$ $0.592\pm$ $0.202$
HM-GNN (+expert) ${0.201\pm}$ 0.0090.009{0.009}0.009 ${0.895\pm}$ 0.0190.019{0.019}0.019 $0.451\pm$ $0.025$ $0.408\pm$ $0.095$ $\mathbf{0.999\pm}$ $\mathbf{0.002}$ $\mathbf{1.000\pm}$ $\mathbf{0.000}$ ${0.681\pm}$ 0.0240.024{0.024}0.024 ${0.587\pm}$ 0.0750.075{0.075}0.075 ${0.185\pm}$ 0.0160.016{0.016}0.016 ${0.926\pm}$ 0.0390.039{0.039}0.039 ${0.345\pm}$ 0.1490.149{0.149}0.149 ${0.547\pm}$ 0.2950.295{0.295}0.295
Ours (-expert) ${0.075\pm}$ 0.0030.003{0.003}0.003 ${0.990\pm}$ 0.0010.001{0.001}0.001 $\mathbf{0.288\pm}$ $\mathbf{0.048}$ ${0.765\pm}$ 0.1460.146{0.146}0.146 ${0.994\pm}$ 0.0010.001{0.001}0.001 ${0.999\pm}$ 0.0000.000{0.000}0.000 ${0.671\pm}$ 0.0200.020{0.020}0.020 ${0.659\pm}$ 0.0470.047{0.047}0.047 ${0.044\pm}$ 0.0150.015{0.015}0.015 ${{0.995\pm}}$ 0.0040.004{0.004}0.004 ${0.268\pm}$ 0.0840.084{0.084}0.084 ${0.738\pm}$ 0.1480.148{0.148}0.148
Ours ${0.045\pm}$ 0.0030.003{0.003}0.003 ${0.996\pm}$ 0.0010.001{0.001}0.001 ${0.295\pm}$ 0.0490.049{0.049}0.049 $\mathbf{0.796\pm}$ $\mathbf{0.105}$ ${0.996\pm}$ 0.0000.000{0.000}0.000 $\mathbf{1.000\pm}$ $\mathbf{0.000}$ ${0.705\pm}$ 0.0070.007{0.007}0.007 $\mathbf{0.711\pm}$ $\mathbf{0.018}$ ${0.028\pm}$ 0.0070.007{0.007}0.007 ${{0.998\pm}}$ 0.0020.002{0.002}0.002 $\mathbf{0.222\pm}$ $\mathbf{0.079}$ $\mathbf{0.819\pm}$ $\mathbf{0.137}$
4.3.2 Comparison with motif-based baselines
@@
To answer question 5), we compare with two baselines. The first, Bag-of-Motifs, ablates our hierarchical information and retains only the motif co-occurrence information. For each molecule, we obtain a feature vector that concatenates a) the occurrence counts of all motifs and b) the Morgan fingerprint of the molecule. We train an XGBoost regressor/classifier on top of these features (details in Appendix E). As shown in Table 4, this baseline has enough capacity to overfit the training data but fails to generalize. This allows us to conclude in the absence of a proper representation, motif occurrence information is not sufficient for generalization. Interestingly, expert-level motifs are not superior to heuristic-based motifs in this featurization. This suggests that the quality of motifs are not relevant in the absence of a hierarchical representation that incorporates the fine-grained features of each individual motif. The second, HM-GNN (Shui & Karypis, 2020), is a SOTA motif-based property predictor that explicitly models motif-molecule and motif-motif relationships using a hetereogenous graph. Furthermore, we endowed the method with our expert motifs since the vanilla version only considers bonds and rings. On both regression datasets, HM-GNN avoids overfitting but does not catch up to our method’s generalization capability. Endowing HM-GNN with our expert motifs enables better fitting of the training data but further hinders generalization. On PTC, HM-GNN is competitive with our method in accuracy but shows a discrepancy in terms of AUC. This is concerning as a lower AUC may imply higher sensitivity to class imbalance (in PTC, there are 45% more negatives than positives) and classification thresholds. Meanwhile, our method can both 1) completely fit the training data ($>0.99$ $R^{2}$, $>99\%$ Acc/AUC), and 2) generalize to the test data, with further regularization potentially leading to even better results. We believe our motif-based representation carries better inductive biases, integrates better with expert motifs, and demonstrates stronger empirical performance.
@@ 4.4.1 Rule extraction from grammar
To answer question 6), we extract context-sensitive grammar rules from our trained model. We perform best-first search over random walk trajectories, beginning with base trajectories corresponding to each group $v\in G$. We only expand trajectories with transition probabilities above a minimum threshold. Each trajectory that reaches a transition with probability of 1 is extracted as a “hard” context-sensitive rule. We depict two such rules in Figure 6, with a more exhaustive compilation in Appendix F.1.